Find the values for m and b in the following system so that the solution to the system is (-3, 4)
5x+7y=b
mx+y=22

Answer by stanbon(48516)

Find the values for m and b in the following system so that the solution to the system is (-3, 4)
5x+7y=b
mx+y=22
-----------
Substitute for x and y so you can solve for m and b:
-15+28=b
b=13
---------
-3m+4=22
-3m=18
m=-6
-------------
cheers,
Stan H.